# Fast SDE-based Monte Carlo dose calculation for proton therapy validated against Geant4

### v1

@@ 3 Sequential proton track SDE
To model the evolution of a proton’s trajectory through the configuration space $\mathcal{C}$, we employ an enhanced version of the SDE first introduced in SDE. We follow the vast majority of physics and nuclear literature, and index our SDE by the inherent ‘track length’ of the sequential proton track that it describes. We define the dynamic evolution of the configuration variables along a sequential proton track via $(Y_{\ell},\ell\geq 0)=((\epsilon_{\ell},r_{\ell},\omega_{\ell}),\ell\geq 0)$, where $\epsilon_{\ell}$ is the energy at track length $\ell\in\mathcal{E}$ of the sequential proton track, $r_{\ell}\in D$ is the position of the sequential proton track at track length $\ell$, and $\omega_{\ell}\in\mathbb{S}_{2}$ is its direction of transport. The process $Y$ represents the random evolution of a sequential proton track in configuration space $\mathcal{C}$}.
@@
Since our focus is on simulating the sequential proton track, we introduce the evolution of $Y$ in terms of its Euler–-Maruyama approximation, a standard numerical scheme for approximately simulating SDE paths. We will work on the lattice track lengths $(\ell=\Delta n,n\geq 0)$, where $\Delta>0$ is a small increment of path length. The collective evolution of $Y_{n}:=(E_{n},R_{n},\Omega_{n}):=(\epsilon_{\Delta n},r_{\Delta n},\Omega_{\Delta n})$, for $n=0,1,2,\dots$ is governed by
$\boxed{\begin{array}[]{rl}E_{n+1}&=(E_{n}-\varsigma_{1}(Y_{n})\Delta+\min\{\max\{\varsigma_{2}(Y_{n})B_{n},-\varsigma_{1}(Y_{n})\Delta\},\varsigma_{1}(Y_{n})\Delta\})(1-u_{n})\\ &\\ R_{n+1}&=R_{n}+\Omega_{n}\Delta\\ &\\ \displaystyle\Omega_{n+1}&=\Omega_{n}+m(Y_{n})\Xi_{n}+D_{n}\end{array}}$ (1)
@@
  • $\Xi_{n}$ is the time-$\Delta$ increment of spherical Brownian motion on $\mathbb{S}_{2}$, simulated via Algorithm 1 of Mijatovic. It models the aggregate effect of small scattering events on the direction of the proton.
@@
In the next section, we present our choices for the functionals $\sigma_{\rm e},\pi_{\rm e},\sigma_{\rm ne},\pi_{\rm ne},\varsigma_{1},\varsigma_{2}$, and $m$. All are informed by the physics literature so that $Y$ replicates the behaviour of a sequential proton track. This makes it possible to apply the SDE model to new materials simply by entering their chemical composition, mean excitation energy, and estimates of the nuclear scattering cross sections in ENDF format without the need to calibrate any free parameters.
4 Choices of SDE parameters
In this section, we describe our choices of parameters for the SDE introduced in Section 3. We will use well-established physics-based models or, when such models are not available, numerical approximations fitted to experimental nuclear data.
4.1 Inelastic Coulomb scattering
We recall from Section 3 that energy losses due to inelastic Coulomb interactions are incorporated into the SDE (1) through a deterministic rate given by $\varsigma_{1}$ and stochastic fluctuations with volatility given by $\varsigma_{2}$. To justify this model choice, we give a brief exposition on the theory of inelastic interactions (see BethandCorrections for further details). First consider the setting of a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a homogeneous medium consisting of a unique element. Each inelastic interaction involves a certain energy transfer $W\in[0,E]$ from the proton to the electrons of an atom in the medium. This energy transfer is completely characterized by the atomic energy-loss differential cross section $f(W,E)\in\mathbb{R}^{+}$, $W\in[0,E]$ and its moments
@@ 4.1 Inelastic Coulomb scattering
If there are $N$ atoms per $\mathrm{cm}^{3}$, then $(N\sigma^{(0)}_{E})^{-1}$ is the mean free path length, $\sigma^{(1)}_{E}/\sigma^{(0)}_{E}$ the mean energy loss in a collision, and $\sigma^{(k)}_{E}/\sigma^{(0)}_{E}$, $k\geq 2$, the $k$-th moment of the energy loss in a collision. Under idealised assumptions that $N$ is sufficiently large and that interactions occur independently, the energy loss per unit track length is well-approximated by $N\sigma^{(1)}_{E}+Z(N\sigma^{(2)}_{E})^{1/2}$, where $Z\sim\mathcal{N}(0,1)$. Indeed, this is just a consequence of the central limit theorem. Thus, for the SDE we choose $\varsigma_{1}(x)=N\sigma^{(1)}_{E}$ and $\varsigma_{2}(x)=(N\sigma^{(2)}_{E})^{1/2}$. Explicit values of $N\sigma^{(1)}_{E}$ and $N\sigma^{(2)}_{E}$ are given by the well-established theory of the Bethe–Bloch formula which is discussed in detail in the following two sections.
4.1.1 The Bethe–Bloch formula for mean energy loss
The current state of the art model for mean energy loss from inelastic collisions is given by the Bethe–Bloch formula, along with its corrections. This quantum theory of stopping, originally due to Bethe, is based on the relativistic plane-wave Born approximation (Bethesimple). The Bethe–Bloch formula without corrections (Gott, Section 3.2 and Appendix D) per unit track length reads
@@ 4.1.1 The Bethe–Bloch formula for mean energy loss
where $\rho$ is the density of the medium in g/cm3, $Z$ is the atomic number of the medium, $A$ is the atomic mass of the medium, $m_{p}$ is the mass of a proton, $m_{e}$ is the mass of an electron, $c$ is the speed of light, and $I$ is the mean excitation energy of the medium. Importantly, this derivation assumes the idealised setting mentioned in the previous section. To account for this, several additional correction terms exist which result in a more accurate model. However, since the scope of this article is modeling the energy deposition of 100–150 MeV protons, we exclude correction terms due to them having little impact in this setting, as justified below.
There are four correction terms to the Bethe–Bloch formula, known as the shell, density, Lindhard–Sørensen, and the Barkas correction term, respectively. Their details can be found in equation Sections V–VII of BethandCorrections. All are known to be negligible in the proton energy range of clinical interest (BISCHEL1992345; MBAGWU2025112354; BethandCorrections; EnergyStragglingSalvat).
To extend (3) to materials consisting of more than one element we use the so-called Bragg-additivity rule. This assumes that the material can be treated as a uniform mix of each of its constituting elements, and that interactions are independent between elements. In this case, the Bethe–Bloch formula reads (Gott, Eq. (8))
@@
Lastly, we note that due to the idealised setting used to derive the Bethe-Bloch formula, equation (3) performs poorly when $E$ is close to 0. Indeed, (3) tends to -$\infty$ as $E\rightarrow 0$. To account for this problem, for the results presented in Section 6, protons are absorbed upon reaching a kinetic energy of 0.05 MeV and their remaining energy is deposited at their current position. The choice of absorption energy threshold ws determined by numerical evaluation of or implementation of (4) to ensure non-negative values, and is low enough that the remaining proton range is below 0.1 mm. Hence, the output of our simulations is unaffected for all practical purposes. An alternative approach using a logarithmic transform of the energy to guarantee non-negative energy for all track length has been considered in chronholm2025geometryenergysensitivitystochastic.
4.1.2 Energy straggling
The original derivation of $N\sigma^{(2)}_{E}$ dates back to Bohr (BohrStraggling, Eq. (3.4.5)). However, this approach neglects the binding of atomic electrons and assumes interactions with non-relativistic free electrons at rest. A more accurate formula based on the relativistic plane-wave Born approximation (EnergyStragglingSalvat; jackson1975classical) reads
@@ 4.1.2 Energy straggling
where $N_{A}$ is Avogadro’s constant, $\gamma$ is the Lorentz factor, $\alpha$ is the fine structure constant, $\hbar$ is the reduced Planck constant, and all other variables are as in Section 4.1.1. Identically to (3), due to the idealised setting assumed when deriving (5), an accurate formula for $\varsigma_{2}$ requires additional correction terms. However, for the clinical energy range of protons with 0–150 MeV, the results of EnergyStragglingSalvat show that they have negligible impact when the atomic number of the material is sufficiently small (less than 40), and between a 1–5% impact for larger atomic numbers. Furthermore, since the effect of energy straggling is small in itself, these errors will be of an order of magnitude smaller when compared to the total energy loss.
@@
4.2 Elastic scattering
We recall from Section 2.1 that elastic scattering is divided into Coulomb and nuclear scattering events, and the interference between them. At small angles, the effect of elastic scattering is dominated by Coulomb contribution (osti_1425114, Section 6.2.7). We recall that the SDE (1) models small angle elastic scattering through a diffusion on the sphere with volatility $m$. Thus, we fit $m$ using a physics-based Coulomb scattering model which is described in detail in Section 4.2.1. The SDE (1) models large angle elastic scattering as a point process with rate $\sigma_{e}$ and jump density $\pi_{e}$. Since no satisfactory theory exists to describe the contributions from nuclear and nuclear-Coulomb elastic scattering, these parameters are fitted from the nuclear data libraries ENDF/B-VIII.1 and JEFF-4.0 which is described in detail in Section 4.2.2.
4.2.1 Molière’s theory for small-angle elastic scattering
The state of the art for modelling Coulomb scattering is given by the theory of Molière (Moliere). The scattering density derived by Molière’s has a Gaussian mode, but tails which are much heavier than Gaussian (Gott). Thus, we model small-angle elastic Coulomb scattering by taking a Gaussian approximation to Moliére’s distribution. Specifically, we use the Lynch–Dahl approximation to Moliére multiple scattering (lynch/dahl:1991, Eq. (7)) which is given as follows. Consider a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ traveling through a medium consisting of a single element. Transforming to spherical coordinates such that $\Omega$ maps to $(1,0,0)$, after a path length $z$ cm, the azimuthal angle is uniform in $0,2\pi]$ and the polar angle is the absolute value of a Gaussian with mean 0 and standard deviation $\sigma_{E}$ given by:
$\chi_{c}^{2}$ $=\frac{0.157Z(Z+1)z}{A\rho(pv)^{2}},$
$\chi_{\alpha}^{2}$ $=2.007\times 10^{-5}Z^{2/3}\frac{1+3.34(Z\alpha/v)^{2}}{(pc)^{2}},$
@@ 4.2.1 Molière’s theory for small-angle elastic scattering
$\sigma_{E}$ $=\sqrt{\frac{\chi_{c}^{2}}{1+F^{2}}\Big(\frac{1+w}{w}\log(1+w)-1\Big)},$
where $p$ is the momentum in $\text{MeV}/c$, $v$ the particles velocity, and all other parameters are as in Section 4.1. Note that the corresponding formula of lynch/dahl:1991 is missing the square root on the final line. Here, $F\in(0,1)$ is a truncation parameter describing the central fraction of the Coulomb scattering distribution taken into account. This is necessary because the tails of the distribution are not integrable, and we take $F=0.98$. This approximation is accurate, and advantageous because it does not use the radiation length of the medium, which is difficult to measure precisely.
@@
$m(x)=\frac{\sigma_{E}}{z^{1/2}}.$
For our simulations in Section 6, we take $z=0.05$ cm and use Algorithm 1 of Mijatovic to compute fast and exact realisations of spherical Brownian motion.
4.2.2 Large angle elastic scattering via experimental data
As shown in Gott, the Lynch–Dahl approximation accurately models the effects of Coulomb elastic scattering for scattering angles up to $\sim 2.5\sigma_{E}$. Beyond that, the light Gaussian tails fall too quickly to match the heavier tails of Molière’s distribution. Computing these heavier tails requires an expensive root finding step (Gott, Eq. (19)). Molière’s theory also gives no insight into the contributions from nuclear and nuclear-Coulomb effects. Thus, we turn instead to nuclear data libraries to model elastic scattering angles above $2.5\sigma_{E}$. Also, we note that although $2.5\sigma_{E}$ is energy-dependent, we fix $2.5\sigma_{E}=0.04$ radians as a cut-off throughout our simulations in Section 6. This is done for the sake of computational efficiency, since a change in $2.5\sigma_{E}$ requires a re-evaluation of the nuclear data. Furthermore, for energies above 10 MeV, 0.04 radians is a good approximation of $2.5\sigma_{E}$ in water.
Consider a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a medium consisting of a single element. We primarily use the ENDF/B-VIII.1 data library, and refer to JEFF-4.0 for elements not included in ENDF/B-VIII.1. For all elements except hydrogen, the scattering cross sections are given by LAW=5 and LTP=12 (osti_1425114, Section 6.2.7). This format expresses the scattering cross section as
@@ 4.2.2 Large angle elastic scattering via experimental data
where $\mu$ is the cosine of the polar scattering angle, $\sigma_{R}^{c}(\mu,E)$ is the differential Coulomb scattering cross section in the center-of-mass frame (conversion to the lab frame is given in Appendix A.1) given by Rutherford’s formula with electronic screening ignored, which reads
@@
with parameters as in Section 4.1, and $\sigma_{NI}^{c},P_{NI}$ given by experimental data to account for nuclear and nuclear-Coulomb effects. Transforming (6) to a scattering rate per cm of path length travelled gives
@@
The integral of $\sigma_{R}^{c}(\mu,E)$ is calculated explicitly. The integral of $P_{NI}(\mu,E))$ is calculated numerically by first fitting a cubic B-spline for $P_{NI}(\mu,E))$, then integrating the resulting spline. The density of the scattering angle is given by
$\pi_{e}(E;\mu)=\frac{(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))}{\int_{-1}^{\cos(2.5\sigma_{E})}(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))\mathrm{d}\mu},\quad\mu\in[-1,\cos(2.5\sigma_{E})].$ (8)
@@
$\Pi_{e}(E;\nu)=\frac{\int_{-1}^{\nu}(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))\mathrm{d}\mu}{\int_{-1}^{\cos(2.5\sigma_{E})}(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))\mathrm{d}\mu},\quad\nu\in[-1,\cos(2.5\sigma_{E})],$ (9)
@@
4.3 Inelastic proton-nucleus collision and scatter
As before, consider a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a medium consisting of a single element. We recall from (1) that inelastic proton-nucleus interactions occur in the SDE at an element-dependent rate $\sigma_{ne}$, and the outgoing energy and angle of the proton after such an interaction is given by the density $\pi_{ne}$. There exists no satisfactory theory to describe the effects of these interactions, so we use experimental data to model $\sigma_{ne}$ and $\pi_{ne}$. As for large elastic scattering, we use the ENDF/B-VIII.1 and JEFF-4.0 nuclear data libraries. These libraries use the Kalbach-Mann systematics representation to model inelastic proton-nucleus interactions (Kalbach_systematics) which is given as follows. The scattering rate per cm of path length travelled, $\sigma_{ne}(E)$, is given for a lattice of points in $\mathcal{E}$ which we fit to a cubic B-spline to extend to $\mathcal{E}$. Furthermore, for a given lattice point $E_{i}\in\mathcal{E}$, the marginal distribution of the outgoing energy, which we denote $\pi_{ne}(E_{i};\mathrm{d}u)$, along with a fitted parameter for the outgoing angle density, denoted $r(E_{i},u)\in[0,1]$, are given for a lattice of points in $\mathcal{E}\setminus(E_{i},\infty]$ which we extend to $\mathcal{E}\setminus(E_{i},\infty]$ and $[0,1]$ respectively using cubic B-splines. Finally, the distribution of the cosine of the polar scattering angle in the center-of-mass frame (conversion to the lab frame is given in Appendix A.1) given incident energy $E$ and outgoing energy $u$ is given by

### v2

@@ 2.1.2 Sequential proton track SDE formulation
To model the evolution of a proton’s trajectory through the configuration space $\mathcal{C}$, we employ an enhanced version of the SDE first introduced in Crossley et al. (2025). We follow the usual physics and nuclear literature convention and index our SDE by the inherent ‘track length’ of the sequential proton track which it describes. We define the dynamic evolution of the configuration variables along a sequential proton track via $(Y_{\ell},\ell\geq 0)=((\epsilon_{\ell},r_{\ell},\omega_{\ell}),\ell\geq 0)$, where $\epsilon_{\ell}$ is the energy at track length $\ell\in\mathcal{E}$ of the sequential proton track, $r_{\ell}\in D$ is the position of the sequential proton track at track length $\ell$, and $\omega_{\ell}\in\mathbb{S}_{2}$ is its direction of transport. The process $Y$ depicts the stochastic evolution of a sequential proton track in configuration space $\mathcal{C}$}.
@@
Since our focus is on simulating the sequential proton track, we introduce the evolution of $Y$ in terms of its Euler–Maruyama approximation: a standard numerical scheme for approximately simulating SDE paths. We will work on the lattice track lengths $(\ell=\Delta n,n\geq 0)$, where $\Delta>0$ is a small increment of path length. The collective evolution of $Y_{n}:=(E_{n},R_{n},\Omega_{n}):=(\epsilon_{\Delta n},r_{\Delta n},\Omega_{\Delta n})$, for $n=0,1,2,\dots$ is governed by
$\boxed{\begin{array}[]{rl}E_{n+1}&=(E_{n}-\varsigma_{1}(Y_{n})\Delta+\min\{\max\{\varsigma_{2}(Y_{n})B_{n},-\varsigma_{1}(Y_{n})\Delta\},\varsigma_{1}(Y_{n})\Delta\})(1-u_{n})\\ &\\ R_{n+1}&=R_{n}+\Omega_{n}\Delta\\ &\\ \displaystyle\Omega_{n+1}&=\Omega_{n}+\Xi_{n}(m(Y_{n}))+D_{n}\end{array}}$ (1)
@@
  • $\Xi_{n}(m(Y_{n}))$ is the time-$\Delta$ increment of spherical Brownian motion on $\mathbb{S}_{2}$ with volatility $m(Y_{n})$, simulated via Algorithm 1 of Mijatović et al. (2020). It models the aggregate effect of small scattering events on the direction of the proton.
@@
In the following sections, we present our choices for the functionals $\sigma_{\rm e},\pi_{\rm e},\sigma_{\rm ne},\pi_{\rm ne},\varsigma_{1},\varsigma_{2}$, and $m$. All are informed by physics so that $Y$ replicates the behaviour of sequential proton track. This makes it possible to apply the SDE model to new materials simply by entering their chemical composition, mean excitation energy, and estimates of the nuclear scattering cross sections in ENDF format without the need to calibrate any free parameters.
2.1.3 Parameterisation for inelastic Coulomb scattering
We recall from Section 2.1.2 that energy losses due to inelastic Coulomb interactions are incorporated into the SDE (1) through two components: a deterministic stopping rate given by $\varsigma_{1}$ and stochastic fluctuations (energy straggling) with volatility given by $\varsigma_{2}$. To justify this model choice, we give a brief exposition on the theory of inelastic interactions (see Salvat (2022) for further details). First consider the setting of a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a homogeneous medium consisting of a unique element. Each inelastic interaction involves an energy transfer $W\in[0,E]$ from the proton to the electrons of an atom in the medium. This energy transfer is completely characterized by the energy-dependent atomic energy-loss differential cross section $f(W,E)\in\mathbb{R}^{+}$, $W\in[0,E]$ and its moments
@@ 2.1.3 Parameterisation for inelastic Coulomb scattering
If there are $N$ atoms per $\mathrm{cm}^{3}$, then $(N\sigma^{(0)}_{E})$ is the mean free path length, $\sigma^{(1)}_{E}/\sigma^{(0)}_{E}$ the mean energy loss in a collision, and $\sigma^{(k)}_{E}/\sigma^{(0)}_{E}$, $k\geq 2$, the $k$-th moment of the energy loss in a collision. Under idealised assumptions that $N$ is sufficiently large and that interactions occur independently, the energy loss per unit track length is well-approximated by $N\sigma^{(1)}_{E}+Z(N\sigma^{(2)}_{E})^{1/2}$, where $Z\sim\mathcal{N}(0,1)$. Indeed, this is just a consequence of the central limit theorem. Thus, for the SDE we choose $\varsigma_{1}(x)=N\sigma^{(1)}_{E}$ and $\varsigma_{2}(x)=(N\sigma^{(2)}_{E})^{1/2}$. Explicit values of $N\sigma^{(1)}_{E}$ and $N\sigma^{(2)}_{E}$ are given by the well-established theory of the Bethe–Bloch formula with energy straggling.
The Bethe–Bloch formula for mean energy loss. The current state of the art model for mean energy loss from inelastic collisions is given by the Bethe–Bloch formula, along with its corrections. This quantum theory of stopping, originally due to Bethe, is based on the relativistic plane-wave Born approximation (Fano, 1963). The Bethe–Bloch formula without corrections (Gottschalk, 2018, Section 3.2 and Appendix D) per unit track length reads
@@
where $\rho$ is the density of the medium in g/cm3, $Z$ is the atomic number of the medium, $A$ is the atomic mass of the medium, $m_{p}$ is the mass of a proton, $m_{e}$ is the mass of an electron, $c$ is the speed of light, and $I$ is the mean excitation energy of the medium. Importantly, this derivation assumes the idealised setting mentioned in the previous section. To account for deviations from the idealised setting, several additional correction terms exist which result in a more accurate model. The four correction terms to the Bethe–Bloch formula are known as the shell, density, Lindhard–Sørensen, and the Barkas correction term, respectively. Their details can be found in equation Sections V–VII of Salvat (2022). All are known to be negligible in the proton energy range of clinical interest (Bischel and Hiraoka, 1992; Mbagwu, 2025; Salvat, 2022; Salvat et al., 2022), and hence we neglect them.
To extend (3) to materials consisting of more than one element we use the so-called Bragg-additivity rule. This assumes that the material can be treated as a uniform mix of each of its constituting elements, and that interactions are independent between elements. In this case, the Bethe–Bloch formula reads (Gottschalk, 2018, Eq. (8))
@@
Lastly, we note that due to the idealised setting used to derive the Bethe-Bloch formula, equation (3) performs poorly when $E$ is close to 0. Indeed, (3) tends to -$\infty$ as $E\rightarrow 0$. To account for this problem, for the results presented in Section 3, protons are absorbed upon reaching a kinetic energy of 0.05 MeV and their remaining energy is deposited at their current position. The choice of absorption energy threshold ws determined by numerical evaluation of or implementation of (4) to ensure non-negative values, and is low enough that the remaining proton range is below 0.1 mm. Hence, the output of our simulations is unaffected for all practical purposes. An alternative approach using a logarithmic transform of the energy to guarantee non-negative energy for all track length has been considered in Chronholm and Pryer (2025).
Energy straggling. The original derivation of $N\sigma^{(2)}_{E}$ dates back to Bohr (Bohr, 1987, Eq. (3.4.5)). It neglects the binding of atomic electrons and assumes interactions with non-relativistic free electrons at rest. A more accurate formula based on the relativistic plane-wave Born approximation (Salvat et al., 2022; Jackson, 1975) reads
@@
where $N_{A}$ is Avogadro’s constant, $\gamma$ is the Lorentz factor, $\alpha$ is the fine structure constant, $\hbar$ is the reduced Planck constant, and all other variables are as in (3). Identically to (3), due to the idealised setting assumed when deriving (5), an accurate formula for $\varsigma_{2}$ requires additional correction terms. However, for the clinical energy range of protons with 0–150 MeV, the results of Salvat et al. (2022) show that they have negligible impact when the atomic number of the material is sufficiently small (less than 40), and between a 1–5% impact for larger atomic numbers. Furthermore, since the effect of energy straggling is small in itself, these errors will be of an order of magnitude smaller when compared to the total energy loss.
@@
2.1.4 Parameterisation for elastic scattering
We recall from Section 2.1.1 that elastic scattering is divided into Coulomb and nuclear scattering events, and the interference between them. At small angles, the effect of elastic scattering is dominated by Coulomb contribution (Trkov and Brown, 2018, Section 6.2.7). We recall that the SDE (1) models small angle elastic scattering through a diffusion on the sphere with volatility $m$. Thus, we fit $m$ using the state of the art for modelling Coulomb scattering, given by the theory of Molière (Bethe, 1953). The SDE (1) models large angle elastic scattering as a point process with rate $\sigma_{e}$ and jump density $\pi_{e}$. Since no satisfactory theory exists to describe the contributions from nuclear and nuclear-Coulomb elastic scattering, these parameters are fitted from the nuclear data libraries ENDF/B-VIII.1 and JEFF-4.0. Both small and large angle contributions are described in detail below.
Molière’s theory for small-angle elastic scattering. The scattering density derived by Molière’s has a Gaussian mode, but tails which are much heavier than Gaussian (Gottschalk, 2018). Thus, we model small-angle elastic Coulomb scattering by taking a Gaussian approximation to Moliére’s distribution. Specifically, we use the Lynch–Dahl approximation to Moliére multiple scattering (Lynch and Dahl, 1991, Eq. (7)) which is given as follows. Consider a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a medium consisting of a single element. Relative to the initial direction of travel $\Omega$, after a path length $z$ cm, the new direction of travel in spherical coordinates has azimuthal angle uniform in $0,2\pi]$ and energy-dependent polar angle that is the absolute value of a Gaussian with mean 0 and standard deviation $\sigma(E)$ given by:
$\chi_{c}^{2}$ $=\frac{0.157Z(Z+1)\rho z}{A(p\beta)^{2}},$
$\chi_{\alpha}^{2}$ $=2.007\times 10^{-5}Z^{2/3}\frac{1+3.34(Z\alpha/\beta)^{2}}{p^{2}},$
@@ 2.1.4 Parameterisation for elastic scattering
$\sigma(E)$ $=\sqrt{\frac{\chi_{c}^{2}}{1+F^{2}}\Big(\frac{1+w}{w}\log(1+w)-1\Big)},$
where $p$ is the momentum in $\text{MeV}/c$, and all other parameters are as in Section 2.1.3. Note that the corresponding formula of Lynch and Dahl (1991) is missing the square root on the final line. Here, $F\in(0,1)$ is a truncation parameter describing the central fraction of the Coulomb scattering distribution taken into account. This is necessary because the tails of the distribution are not integrable, and thus we take $F=0.98$. This approximation is accurate, and advantageous because it does not use the radiation length of the medium, which is difficult to measure precisely.
@@
$m(x)=\frac{\sigma(E)}{z^{1/2}}.$
For our simulations in Section 3, we take $z=0.05$ cm and use Algorithm 1 of Mijatović et al. (2020) to compute fast and exact realisations of spherical Brownian motion.
Large angle elastic scattering via experimental data. As shown in Gottschalk (2018), the Lynch–Dahl approximation accurately models the effects of Coulomb elastic scattering for scattering angles up to $\sim 2.5\sigma(E)$. Beyond that, the light Gaussian tails fall too quickly to match the heavier tails of Molière’s distribution. Computing these heavier tails requires an expensive root finding step (Gottschalk, 2018, Eq. (19)). Molière’s theory also gives no insight into the contributions from nuclear and nuclear-Coulomb effects. Thus, we turn instead to nuclear data libraries to model elastic scattering angles above $2.5\sigma(E)$. Also, we note that although $2.5\sigma(E)$ is energy-dependent, we fix $2.5\sigma(E)=0.02$ radians as a cut-off throughout our simulations in Section 3. This is done for the sake of computational efficiency, since a change in $2.5\sigma(E)$ requires a re-evaluation of the nuclear data. Furthermore, for energies above 10 MeV, 0.02 radians is a good approximation of $2.5\sigma(E)$ for elements seen in proton beam therapy.
We primarily use the ENDF/B-VIII.1 data library, and refer to JEFF-4.0 for elements not included in ENDF/B-VIII.1. For all elements except hydrogen, the scattering cross sections are given by LAW=5 and LTP=12 (Trkov and Brown, 2018, Section 6.2.7). Consider a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a medium consisting of a single element. This format expresses the scattering cross section as
@@
where $\mu$ is the cosine of the polar scattering angle, $\sigma_{R}^{c}(\mu,E)$ is the differential Coulomb scattering cross section in the center-of-mass frame given by Rutherford’s formula with electronic screening ignored (conversion to the lab frame is given in Appendix A.1), which reads
@@
with parameters as in Section 2.1.3, and $\sigma_{NI}^{c},P_{NI}$ given by experimental data to account for nuclear and nuclear-Coulomb effects. Transforming (6) to a scattering rate per cm of path length travelled gives
@@
The integral of $\sigma_{R}^{c}(\mu,E)$ is calculated explicitly. The integral of $P_{NI}(\mu,E)$ is calculated numerically by first fitting a cubic B-spline for $P_{NI}(\mu,E)$, then integrating the resulting spline. The density of the scattering angle is given by
$\pi_{e}(E;\mu)=\frac{(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))}{\int_{-1}^{\cos(2.5\sigma_{E})}(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))\mathrm{d}\mu},\quad\mu\in[-1,\cos(2.5\sigma(E))].$ (8)
@@
$\Pi_{e}(E;\nu)=\frac{\int_{-1}^{\nu}(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))\mathrm{d}\mu}{\int_{-1}^{\cos(2.5\sigma_{E})}(\sigma_{R}^{c}(\mu,E)+\sigma_{NI}^{c}(E)P_{NI}(\mu,E))\mathrm{d}\mu},\quad\nu\in[-1,\cos(2.5\sigma(E))],$ (9)
@@
2.1.5 Parameterisation for inelastic proton-nucleus interactions
As before, consider a proton with configuration $x=(E,R,\Omega)\in\mathcal{C}$ travelling through a medium consisting of a single element. We recall from (1) that inelastic proton-nucleus interactions occur in the SDE at an element-dependent rate $\sigma_{ne}$, and the outgoing energy and angle of the proton after such an interaction is given by the density $\pi_{ne}$. There exists no satisfactory theory to describe the effects of these interactions, so we use experimental data to model $\sigma_{ne}$ and $\pi_{ne}$. As for large elastic scattering, we use the ENDF/B-VIII.1 and JEFF-4.0 nuclear data libraries. These libraries use the Kalbach-Mann systematics representation to model inelastic proton-nucleus interactions (Kalbach, 1988) which is given as follows. The scattering rate per cm of path length travelled, $\sigma_{ne}(E)$, is given for a lattice of points in $\mathcal{E}$ which we fit to a cubic B-spline to extend to $\mathcal{E}$. For $E\in\mathcal{E}$, $u\in[0,1]$, let $\pi_{ne}(E;\mathrm{d}u)$ denote the marginal distribution of the outgoing energy given incident energy $E$. For a given set of lattice points $E_{i}\in\mathcal{E}$, $U_{j}\in[0,1]$, $\pi_{ne}(E_{i};U_{j})$ is given along with a fitted parameter for the outgoing angle density, denoted $r(E_{i},U_{j})\in[0,1]$. We extend these to $\mathcal{E}$ and $[0,1]$ respectively using linear interpolation. Finally, the distribution of the cosine of the polar scattering angle in the center-of-mass frame (conversion to the lab frame is given in Appendix A.1) given incident energy $E$ and outgoing energy (1u)E(1-u)E is given by